# Quantum gravity cannot be both consistent and complete

### v1

@@ Abstract
General relativity, despite its profound successes, fails as a complete theory due to presence of singularities. While it is widely believed that quantum gravity has the potential to be a complete theory, in which spacetime consistently emerges from quantum degrees of freedom through computational algorithms, we demonstrate that this goal is fundamentally unattainable. Gödel’s theorems establish that no theory based on computational algorithms can be both complete and consistent, while Tarski’s undefinability theorem demonstrates that even within quantum gravity, or any computational framework, a fully consistent internal determination of true propositions is impossible. Chaitin’s incompleteness theorem further reinforces this conclusion, revealing intrinsic limits to any computational theory. We discuss some possible consequences for descriptions of physical systems, and note that a non-algorithmic approach is essential for any theory of everything.
@@ I Introduction
These developments strongly align with Wheeler’s ”it from bit” hypothesis, along with its contemporary formulations within string theory Jafferis:2022crx ; VanRaamsdonk:2020ydg and loop quantum LQG Makela:2019vgf , which assert that information fundamentally underpins physical reality Wheeler1990 . The inability of classical theories to describe singularities further supports this perspective: spacetime, as an emergent phenomenon, cannot account for regions where its underlying quantum information structure cannot be described in terms of spacetime geometry. While ‘it’–representing spacetime and matter fields in it–is not complete, it is expected that ‘bit’–representing quantum gravity–could be complete and consistent computational ‘theory of everything’. We will argue here, however, that it is not possible to construct such a theory based on algorithmic computations.
@@ II Quantum Gravity
Quantum gravity aims to unify the principles of quantum mechanics and general relativity within a single consistent framework. Despite significant advancements in approaches such as string theory Green1987 ; Polchinski1998 and loop quantum gravity (LQG) Rovelli2004 ; Thiemann2007 , a complete and definitive theory remains elusive. So, we have different approaches to quantum gravity, and based on those different approaches multiple axiomatic constructions of quantum gravity have been proposed Witten:1985cc ; Ziaeepour:2021ubo ; Faizal2024 ; bombelli1987spacetime ; Majid:2017bul ; DAriano:2016njq ; Arsiwalla:2021eao . In general, we first note that any viable candidate for quantum gravity must adhere to fundamental axioms that ensure both physical and mathematical coherence. Furthermore, such a formal mathematical axiomatic system, $\mathcal{F}_{QG}$, generates spacetime rather than existing within it Faizal2023 . The essential axioms required for quantum gravity to qualify as a theory of everything, can be summarized as follows:
@@
4. Completeness: The theory must describe phenomena across all scales, from the Planck scale to cosmological scales. This includes the resolution of singularities, such as those found in black holes and the big bang Bojowald2001 ; Mathur2005 . A complete theory should also encompass the origin of universe, which is attempted via quantum cosmology in LQG Ashtekar2006 ; Rovelli2004 or colliding branes in string theory Khoury:2001wf . It should be consistent with all observational data.
@@ III Gödel’s and Tarski’s Theorems in Quantum Gravity
In this section, we demonstrate how Gödel’s incompleteness theorems Godel1931 ; Smith2007 and Tarski’s undefinability theorem Tarski1933 ; Tarski1983 apply to quantum gravity if it is formulated as a formal axiomatic system Faizal2024 ; Faizal20240 . This formal system $\mathcal{F}_{QG}$, apart from consistency ($\mathcal{F}_{QG}$ does not derive any contradictions, i.e., $\mathcal{F}_{QG}\not\vdash\bot$.) will only be required to be sufficiently expressive and effectively axiomatizable for these theorems to hold.
As $\mathcal{F}_{QG}$ is consistent and sufficiently expressive, then there exists a scientific fact/statement $G$ in the language of $\mathcal{F}_{QG}$ such that: $\mathcal{F}_{QG}\not\vdash G\quad$and $\mathcal{F}_{QG}\not\vdash\neg G$ Godel1931 ; Smith2007 . Assign a unique Gödel number $\text{GN}(\phi)\in\mathbb{N}$ to every scientific fact/statement $\phi$ in $\mathcal{F}_{QG}$. This Gödelization process ensures that such scientific statements in the formal system can be encoded as integers. Now we can define the provability predicate $\text{Prov}_{\mathcal{F}_{QG}}(x)$, which asserts that $x$ is the Gödel number of a provable statement in $\mathcal{F}_{QG}$. We can also construct a self-referential statement $G$ such that: $G\equiv\neg\text{Prov}_{\mathcal{F}_{QG}}(\text{GN}(G)).$ This statement asserts, “$G$ is not provable in $\mathcal{F}_{QG}$.” Here, we analyze the two cases: If $\mathcal{F}_{QG}\vdash G$, then $\text{Prov}_{\mathcal{F}_{QG}}(\text{GN}(G))$ holds, contradicting the definition of $G$. However, if $\mathcal{F}_{QG}\vdash\neg G$, then $\mathcal{F}_{QG}$ proves $G$ is not provable. However, this implies $G$ is true, leading to a contradiction. Thus, $G$ is undecidable within $\mathcal{F}_{QG}$.
The undecidability of $G$ implies that there are scientific facts/statements in a purely algorithmic axiomatic system such as any purported theory of quantum gravity that are true by the very construction of the system, but inherently not provable within $\mathcal{F}_{QG}$. In fact, it has already been proposed that the measurement problem in quantum mechanics is related to Gödel theorem in objective collapse models Penrose1996 ; Diosi1987 .
Furthermore, while $\mathcal{F}_{QG}$ is consistent, $\mathcal{F}_{QG}$ cannot prove its own consistency: $\mathcal{F}_{QG}\not\vdash\text{Con}(\mathcal{F}_{QG}),$ where $\text{Con}(\mathcal{F}_{QG})$ is the statement: $\text{Con}(\mathcal{F}_{QG})\equiv\neg\text{Prov}_{\mathcal{F}_{QG}}(\text{GN}(\bot))$ Godel1931 ; Smith2007 . Here, we assume $\mathcal{F}_{QG}\vdash\text{Con}(\mathcal{F}_{QG})$. By Gödel’s first theorem, there exists a statement $G$ undecidable in $\mathcal{F}_{QG}$. If $\mathcal{F}_{QG}\vdash\text{Con}(\mathcal{F}_{QG})$, then $\mathcal{F}_{QG}$ can prove $G$, contradicting the undecidability of $G$. Thus, $\mathcal{F}_{QG}\not\vdash\text{Con}(\mathcal{F}_{QG})$. The consistency of $\mathcal{F}_{QG}$ must be established externally, through a meta-theoretical framework, which is not algorithmic.
@@
Now we will analyze the implication of Tarski’s undefinability theorem for $\mathcal{F}_{QG}$ Faizal20241 . The truth predicate $\text{True}_{\mathcal{F}_{QG}}(x)$ for statements in $\mathcal{F}_{QG}$ cannot be defined within $\mathcal{F}_{QG}$ Tarski1933 ; Tarski1983 . To demonstrate this, assume $\text{True}_{\mathcal{F}_{QG}}(x)$ is definable within $\mathcal{F}_{QG}$. Then construct a statement $\psi$ such that: $\psi\equiv\neg\text{True}_{\mathcal{F}_{QG}}(\text{GN}(\psi)).$ By the definition of $\text{True}_{\mathcal{F}_{QG}}(x)$, $\psi$ is true if and only if $\psi$ is not true, which is a contradiction. The undefinability of the truth predicate suggests that the “truth” of certain scientific statements in any algorithmic proposed theory such as quantum gravity, again requires external meta-theoretical frameworks for full comprehension. It is not possible to fully internally define scientific truth computationally. It had been argued by Weinberg that if we had a final computational theory of everything, a full theory of quantum gravity, then all scientific truths would be computationally derived from the axioms of that theory weinberg2011dreams . Tarski’s undefinability theorem shows that this is impossible.
@@ IV Chaitin’s Incompleteness Theorem
Even though formal mathematical systems are based on logic, and Gödel’s and Tarski’s theorems demonstrate implicitly the limitations of quantum gravity, real calculations in quantum gravity are done using computational algorithms. If we want to explicitly analyze such limitations, we need to analyze similar limitations in algorithmic information theory. Chaitin’s incompleteness theorem chaitin1975theory ; Chaitin2004 is a profound result in algorithmic information theory, which results in similar fundamental limitations as Gödel’s theorems (and is directly related to them kritchman2010surprise ). In essence, it states that in any consistent and sufficiently expressive formal system, there exists a bound on the provability of statements based on their algorithmic complexity. This theorem complements Gödel’s incompleteness theorems by demonstrating that the inability to prove certain truths arises from their informational complexity.
We now explore how Chaitin’s theorem applies to quantum gravity, formulated as a formal system. As $\mathcal{F}_{QG}$ denotes the formal system describing quantum gravity, with axioms $\{A_{1},A_{2},\dots\}$. Chaitin’s theorem states that there exists a constant $K_{\mathcal{F}_{QG}}$, determined by the axioms of $\mathcal{F}_{QG}$, such that no statement $S$ with Kolmogorov complexity $K(S)>K_{\mathcal{F}_{QG}}$ can be proven within $\mathcal{F}_{QG}$. The Kolmogorov complexity $K(S)$ of a statement $S$ is defined as: $K(S)=\min\{|P|:U(P)=S\},$ where $P$ is a program for a universal Turing machine $U$, and $|P|$ denotes the length of $P$ in bits. Now each statement $S$ in $\mathcal{F}_{QG}$ can be encoded as a finite binary string. Let $\text{GN}(S)$ denote the Gödel number of $S$, which allows us to measure the Kolmogorov complexity $K(S)$.
Since $\mathcal{F}_{QG}$ is effectively axiomatizable, all proofs in $\mathcal{F}_{QG}$ can be enumerated as $P_{1},P_{2},\dots$. Let $P_{i}$ be a proof of length $|P_{i}|$, which corresponds to a provable statement $S_{i}$. There are at most $2^{n}$ proofs of length $\leq n$, and thus only finitely many statements $S_{i}$ with $K(S_{i})\leq n$. By the Kraft inequality goldfeld_kraft_proof , the total number of binary strings with Kolmogorov complexity $K(S)\leq n$ is bounded by $12^{n+1}-$ 1. However, the number of binary strings of length $n$ is $2^{n}$, meaning that there exist strings $S$ with $K(S)>n$. These strings cannot be generated by programs of length $\leq n$, and hence cannot correspond to provable statements in $\mathcal{F}_{QG}$. Suppose, for example, that $\mathcal{F}_{QG}$ proves a statement $S$ with $K(S)>K_{\mathcal{F}_{QG}}$. Then there exists a proof $P$ such that $|P|=K(S)>K_{\mathcal{F}_{QG}}$, contradicting the definition of $K_{\mathcal{F}_{QG}}$. Thus, no such statement $S$ can be proven within $\mathcal{F}_{QG}$.
@@
Chaitin’s theorem has profound implications for quantum gravity. Many scientific facts/statements in quantum gravity, such as those describing the microstates of black holes, involve high complexity. The Bekenstein-Hawking entropy, for instance, scales as $S_{BH}\sim A/(4\ell_{p}^{2})$, where $A$ is the horizon area. If information about individual microstates require Kolmogorov complexity exceeding $K_{\mathcal{F}_{QG}}$, they will be unprovable within the formal system of quantum gravity. It has already been argued that quantum gravity prevents its own measurement near the Planck scale Pourhassan:2023jbs ; Pourhassan2025 . Similarly, detailed descriptions of spacetime topology at the Planck scale might involve complexities beyond the provability threshold of $\mathcal{F}_{QG}$, suggesting that some truths in quantum gravity, though valid, would lie outside the reach of computational derivation.

### v3

# Can quantum gravity be both consistent and complete?

@@ Abstract
General relativity, despite its profound successes, fails as a complete theory due to presence of singularities. While it is widely believed that quantum gravity has the potential to be a complete theory, in which spacetime consistently emerges from quantum degrees of freedom through computational algorithms, we argue that this goal could be fundamentally unattainable. We examine how this limitation could emerge in various contexts, depending on whether or not every mathematically valid result is physically realized. In the first case, Gödel’s incompleteness theorems, along with related results by Tarski and Chaitin, imply that no theory formulated as a formal axiomatic system can be complete, and that within any computational framework, a fully consistent internal truth predicate is impossible. In the second case, if only a subset of mathematical truths is realized in nature, we argue that this selection cannot be determined by any purely computational process. Hence, a meta-theoretical approach based on non-algorithmic understanding is indispensable in every case. We discuss some possible consequences of this observation for describing physical systems and note that a non-algorithmic approach should be essential for any theory of everything.
@@ I Introduction
These developments strongly align with Wheeler’s ”it from bit” proposal, along with its contemporary formulations within string theory Jafferis:2022crx ; VanRaamsdonk:2020ydg and loop quantum gravity (LQG) Makela:2019vgf , which assert that information fundamentally underpins physical reality Wheeler1990 . The inability of classical theories to describe singularities further supports this perspective: spacetime, as an emergent phenomenon, cannot account for regions where its underlying quantum information structure cannot be described in terms of spacetime geometry. While ‘it’–representing spacetime and matter fields in it–is not complete, it is expected that ‘bit’–representing quantum gravity–could be complete and consistent computational ‘theory of everything’. We will argue here, however, that it is not possible to construct such a theory based on algorithmic computations.
Here we clarify two distinct philosophical attitudes toward the relation between mathematics, the theory of everything, and physical reality. The first approach can be referred to as Universal Actualism. In this first stance, championed by Weinberg weinberg2011dreams , every mathematically sound deduction from the theory of everything is, in principle, physically actualized. The theory of everything is identified with a maximally expressive set of axioms; once the axioms are fixed, all their theorems are regarded as concrete physical facts about the cosmos. Existence is therefore equated with formal derivability, and the distinction between “possible” and “actual” collapses into pure mathematics. Recent proponents of this view include “Mathematical Universe Hypothesis,” which similarly equates physical existence with membership in the set of consistent mathematical structures tegmark2008mu .
The second philosophical position may be referred to as Selective Actualism. In this approach proposed by Schmidhuber, it is argued that while all computable universes exist as valid abstract programs, only a select subset are physically actualized schmidhuber1997computer . A clear distinction is made between what is a mathematically valid deduction from the theory of everything and what is physically realized—the latter being only a subset of the former. This selection is not determined by computation alone, but rather by a meta-mathematical principle that relies on non-algorithmic understanding.
Although Selective Actualism clearly requires the existence of non-algorithmic understanding, the purpose of this essay is to argue that non-algorithmic understanding should be also essential even within the framework of Universal Actualism. Gödel’s incompleteness theorems, together with related results by Tarski and Chaitin, suggest that no theory expressed as a formal axiomatic system can be both complete and consistent. This suggests that a form of non algorithmic understanding is required, even in a context where all mathematically valid results are expected to be physically realized.
@@ II Quantum Gravity
Quantum gravity aims to unify the principles of quantum mechanics and general relativity within a single consistent framework. Despite significant advancements in approaches such as string theory Green1987 ; Polchinski1998 and loop quantum gravity (LQG) Rovelli2004 ; Thiemann2007 , a complete and definitive theory remains elusive. So, we have different approaches to quantum gravity, and based on those different approaches multiple axiomatic constructions of quantum gravity have been proposed Witten:1985cc ; Ziaeepour:2021ubo ; Faizal2024 ; bombelli1987spacetime ; Majid:2017bul ; DAriano:2016njq ; Arsiwalla:2021eao . In general, we first note that any viable candidate for quantum gravity must adhere to fundamental axioms that ensure both physical and mathematical coherence. Furthermore, such a formal mathematical axiomatic system, $\mathcal{F}_{QG}$, generates spacetime rather than existing within it Faizal2023 . These include both physically motivated axioms of quantum gravity denoted by $\mathcal{F}_{PQG}$, and the axioms of the mathematics quantum gravity is built upon (such as Zermelo-Fraenkel set theory with the Axiom of Choice) denoted by $\mathcal{F}_{MQG}$. Here, we have $\mathcal{F}_{QG}=\mathcal{F}_{PQG}\cup\mathcal{F}_{MQG}$.
The essential axioms required for quantum gravity to qualify as a theory of everything, can be summarized as follows:
@@
4. Completeness  [^6] In this context, completeness should be understood as a complete description of all physical phenomena. It should not be confused with mathematical completeness, which is considered here only in the context of Universal Actualism. In Selective Actualism, completeness means that everything physical should be a computational or Gödelian consequence of the theory, although not every mathematical consequence of the theory is necessarily physical.: The theory must describe phenomena across all scales, from the Planck scale to cosmological scales. This includes the resolution of singularities, such as those found in black holes and the big bang Bojowald2001 ; Mathur2005 . A complete theory should also encompass the origin of universe, which is attempted via quantum cosmology in LQG Ashtekar2006 ; Rovelli2004 or colliding branes in string theory Khoury:2001wf . It should be consistent with all observational data.
@@ III Gödel’s and Tarski’s Theorems in Quantum Gravity
We will begin by distinguishing between the computational consequences and the Gödelian consequences of an axiomatic system describing quantum gravity Godel1931 ; Smith2007 . In any sufficiently expressive, consistent, and effectively axiomatizable formulation of quantum gravity, it is possible to derive certain consequences through computational algorithms. However, these computational outcomes do not encompass the full set of consequences that can be derived from the system.
To begin, we consider a formal system for quantum gravity, which we denote as $\mathcal{F}_{QG}$. This system is assumed to be consistent, meaning that it does not derive any contradictions, but it is also sufficiently expressive and axiomatizable for these theorems to hold. Gödel’s incompleteness theorems state that in any formal system that is sufficiently expressive, there exist mathematical statements that are true within the system but cannot be proven by the system itself Godel1931 ; Smith2007 . In the context of quantum gravity, as the formal system $\mathcal{F}_{QG}$ is consistent and sufficiently expressive, there must be certain mathematical statements, like $G$, which cannot be proved within the system. Such statements are true by virtue of the structure of the system but are not provable using its axioms and rules.
So, certainly, $\mathcal{F}_{QG}$ cannot be both consistent and complete, and this is a direct implication of Gödel’s theorems. The truths that will be true will not be limited to the computational consequences of $\mathcal{F}_{QG}$, but will also include Gödelian consequences of $\mathcal{F}_{QG}$. In Universal Actualism, these Gödelian consequences will definitely correspond to some physical aspects of the theory. Thus, undefinability will be an intrinsic feature of Universal Actualism, and we will require a meta-theoretical approach to address them. On the other hand, in Selective Actualism, two possibilities arise: either some Gödelian consequences are part of the physically important consequences, or none of the Gödelian consequences are expressed as physically important consequences. In the first case, we still have some Gödelian consequences, which are a part of the physical theory and hence need a meta-theoretical approach. The other case requires the identification and elimination of all Gödelian consequences of the theory, and this selection process cannot be algorithmic. So, as this elimination of all the Gödelian consequences is, in itself, non-algorithmic and beyond computation, it again requires a meta-theoretical approach. In any case, we do require a meta-theoretical approach based on a non-algorithmic understanding to construct quantum gravity Faizal2024 ; Faizal20240 .
Additionally, while the system $\mathcal{F}_{QG}$ can be assumed to be consistent, it cannot prove its own consistency. This is a direct consequence of Gödel’s second incompleteness theorem, which asserts that no sufficiently powerful formal system can prove its own consistency. Therefore, $\mathcal{F}_{QG}$ cannot prove the mathematical statement $\text{Con}(\mathcal{F}_{QG})$, which asserts that $\mathcal{F}_{QG}$ is free from contradictions. If $\mathcal{F}_{QG}$ were able to prove its own consistency, it would also be able to prove the mathematical statement $G$, which contradicts the undecidability of $G$. Consequently, the consistency of $\mathcal{F}_{QG}$ must be established externally, relying on a meta-theoretical framework that transcends the formal system itself Faizal2024 ; Faizal20240 . This holds true for both Selective Actualism and Universal Actualism. In Selective Actualism, this is trivially true, as the very definition of the approach does not imply that a fully complete and consistent physical theory can be derived solely from computations based on the axioms of quantum gravity. In Universal Actualism, it holds true due to Gödel’s incompleteness theorems.
We have not yet discussed whether these consequences will have important physical implications or will be trivially true due to the very structure of $\mathcal{F}_{QG}=\mathcal{F}_{PQG}\cup\mathcal{F}_{MQG}$. Here, we would like to note that, on physical grounds, some Gödelian statements of $\mathcal{F}_{QG}$ should have important physical consequences. Different approaches to quantum gravity are based on the quantization of some underlying degrees of freedom and thus obey quantum logic. For example, string theory is based on the quantum theory of strings Green1984 ; Polchinski1998 , while LQG is based on a quantum theory of certain discrete degrees of freedom Ashtekar1986 ; Rovelli2004 . Hence, all approaches to quantum gravity obey quantum logic. However, quantum logic has been demonstrated to be fundamentally undecidable and Gödelian Fritz2021 ; Eisert2012 ; vandenNest2008measurement ; lloyd1993quantum . Even when quantum mechanics has been proposed to be modified by quantum gravity, such as in gravitational collapse models, it has been suggested that the objective collapse in the measurement problem of quantum mechanics itself is related to Gödel’s incompleteness theorem Penrose1996 ; Diosi1987 . Thus, there are fundamental physical consequences of every quantum system that cannot be fully captured by any algorithmic computation. Certainly, these would be present in any formalism of quantum gravity, which would be based on some quantum logic. More directly, aligning with Wheeler’s ‘it from bit’ Wheeler1990 —along with its contemporary formulations within string theory Jafferis:2022crx ; VanRaamsdonk:2020ydg and LQG Makela:2019vgf —we observe that quantum information in quantum gravity constitutes the basis of observable physical phenomena. As quantum logic for this quantum information describing quantum gravity is fundamentally undecidable and Gödelian in nature, certain physical phenomena that follow from the quantum logic of quantum gravity must likewise be undecidable and Gödelian. So, as ‘bit’—or more precisely, ‘qubit’—is certainly Gödelian Fritz2021 ; Eisert2012 ; vandenNest2008measurement ; lloyd1993quantum , ‘it’ must be Gödelian as well.
Furthermore, it has been argued that even the selection of observables in general relativity is Gödelian and thus undecidable Panagiotopoulos2023 . It seems very unlikely that quantum gravitational arguments could make it decidable. In fact, we would possibly calculate them as expectation values of some underlying degrees of freedom. The details of the calculation will depend on the specific approach taken. However, those calculations would be undecidable and Gödelian, as we already know their expectation value has to coincide with the results in general relativity due to the Ehrenfest’s theorem.
@@
We now note that, within the framework of Universal Actualism, it has been argued using Weinberg’s approach that if we had a final computational theory of everything, a complete theory of quantum gravity, then all scientific mathematical truths would be computationally derived from the axioms of that theory weinberg2011dreams . Tarski’s undefinability theorem shows that this is impossible Tarski1933 ; Tarski1983 . Thus, we now turn to the implications of Tarski’s undefinability theorem for $\mathcal{F}_{QG}$. According to Tarski’s theorem, the truth predicate for mathematical statements within $\mathcal{F}_{QG}$ cannot be defined within the system itself. If we assume that the truth predicate, $\text{True}_{\mathcal{F}_{QG}}(x)$, can be defined within $\mathcal{F}_{QG}$, we can construct a self-referential mathematical statement $\psi$ that asserts, ”The statement $\psi$ is not true.” This leads to a paradox, as $\psi$ would be true if and only if it is not true, thus violating the principle of non-contradiction. This contradiction shows that the truth predicate cannot be defined within the formal system.
The undefinability of the truth predicate has profound implications for the understanding of quantum gravity. It suggests that certain truths within the system cannot be fully understood within the system itself and require an external, meta-theoretical framework for their interpretation. This meta-theoretical framework is based on an external truth predicate, which is constructed using non-algorithmic understanding. It forms the deepest aspect of the workings of quantum gravity. So, Tarski’s undefinability theorem reinforces the idea that quantum gravity, like any formal system that seeks to capture fundamental physical reality, cannot be self-contained and requires external structures to fully comprehend its truths Faizal20241 . Selective Actualism is already based on such non-algorithmic, non-computational external structures, which selectively express some consequences of the theory physically. Here, we have argued that this is also true for Universal Actualism, and that the Weinberg proposal is impossible to attain due to Tarski’s undefinability theorem.
Gödel’s incompleteness theorems and Tarski’s undefinability theorem highlight the limitations of any formal, algorithmic approach to understanding quantum gravity. These theorems suggest that quantum gravity cannot be fully captured by an internal, self-contained formal system. Instead, it requires an external meta-theoretical framework to address fundamental questions about provability, truth, and consistency. This underscores the need for an approach based on a non-algorithmic understanding to discuss quantum gravity—one that goes beyond the computational consequences of formal systems and incorporates external truth predicates for a complete description.
@@ IV Chaitin’s Incompleteness Theorem
Even though formal mathematical systems are based on logic, and Gödel’s and Tarski’s theorems demonstrate implicitly the limitations of a theory of quantum gravity purely formulated as a formal axiomatic system, practical calculations in quantum gravity are typically carried out using computational algorithms. To explicitly investigate these limitations, one must therefore examine the constraints within the framework of algorithmic information theory.
@@
Chaitin’s incompleteness theorem chaitin1975theory ; Chaitin2004 is a profound result in algorithmic information theory, which results in similar fundamental limitations as Gödel’s theorems (and is directly related to them kritchman2010surprise ). In essence, it states that in any consistent and sufficiently expressive formal system, there exists a bound on the provability of mathematical statements based on their algorithmic complexity. This theorem complements Gödel’s incompleteness theorems by demonstrating that the inability to prove certain truths arises from their informational complexity. It should be noted that, although algorithmic information theory demonstrates that any formal system contains infinitely many undecidable mathematical statements, only a vanishingly small fraction of such unprovable statements could plausibly relate to specific physical properties. Establishing such a correspondence requires explicit constructions, as shown for instance for specific systems paterek2010 ; purcell2024 .
We will first analyze the mathematical consequences of Chaitin’s theorem for any axiomatic formalization of quantum gravity Witten:1985cc ; Ziaeepour:2021ubo ; Faizal2024 ; bombelli1987spacetime ; Majid:2017bul ; DAriano:2016njq ; Arsiwalla:2021eao , and then discuss some physical implications. For any formal axiomatic system, Chaitin’s theorem applies directly. Accordingly, we can state that for any formal system $\mathcal{F}_{QG}=\mathcal{F}_{PQG}\cup\mathcal{F}_{MQG}$, there exists a constant $K_{\mathcal{F}_{QG}}$, determined by the axioms of $\mathcal{F}_{QG}$, such that no mathematical statement $S$ with Kolmogorov complexity $K(S)$ greater than $K_{\mathcal{F}_{QG}}$ can be proven within the system. The Kolmogorov complexity $K(S)$, is defined as the length of the shortest possible program that can compute $S$ on a universal Turing machine. In other words, it is the minimal number of bits required to describe $S$ algorithmically. Thus, each mathematical statement within $\mathcal{F}_{QG}$ can be encoded as a finite binary string, allowing us to apply Chaitin’s theorem to the formal structure of the theory. However, it remains unclear whether these mathematical limitations are merely a trivial consequence of the fact that $\mathcal{F}_{QG}=\mathcal{F}_{PQG}\cup\mathcal{F}_{MQG}$, or whether they have meaningful physical implications.
It has recently been demonstrated that, due to Chaitin’s theorem, the spacetime foam (i.e., Planckian quantum gravitational fluctuations of spacetime) would become undecidable Faizal2025 . It has also been argued that these fluctuations of spacetime can have detectable signatures on gravitational waves. So, detailed descriptions of spacetime topology at the Planck scale involve complexities beyond the provability threshold of $\mathcal{F}_{QG}$, thus showing that some truths in quantum gravity, though valid, would lie outside the reach of computational derivation. This is a direct consequence of the fact that the spacetime foam is undecidable in quantum gravity as a result of Chaitin’s theorem Faizal2025 .
@@
Furthermore, any systems in quantum gravity, such as those describing the microstates of black holes, involve high complexity. The Bekenstein-Hawking entropy, for instance, scales as $S_{BH}\sim A/(4\ell_{p}^{2})$, where $A$ is the horizon area. The scaling of entropy with the black hole’s surface area suggests that the number of possible microstates grows exponentially with the area, indicating a highly complex structure underlying the microstates of a black hole Balasubramanian:2024rek . If information about individual microstates requires a Kolmogorov complexity exceeding $K_{\mathcal{F}_{QG}}$, they will be unprovable within the formal system of quantum gravity. It has already been argued that quantum gravity prevents its own measurement near the Planck scale Pourhassan:2023jbs ; Pourhassan2025 . In fact, these works argue that this is directly related to incompleteness theorems.
It has already been demonstrated that the Heisenberg uncertainty principle in quantum mechanics can be directly related to Chaitin’s theorem Calude2005 ; Calude2010 . Now, it is possible to study the uncertainty principle for any quantum gravitational system, and it can directly affect the properties of spacetime. The uncertainty principle has been studied in string theory, where it has direct implications for black holes and D-branes Yoneya2000 . Similarly, the uncertainty principle has also been studied in the context of LQG Perlov2018 . The uncertainty principle in quantum gravity has direct implications for quantum cosmology Ohkuwa:2012cm ; Ohkuwa:2012wz . These results can be seen as direct consequences of Chaitin’s theorem on quantum gravity.
Thus, Chaitin’s theorem suggests a fundamental limitation in the axiomatic formalization of quantum gravity. It implies that there exist true mathematical statements in quantum gravity, which have a Kolmogorov complexity beyond the capacity of the formal system to prove. In other words, there are mathematical truths in quantum gravity that cannot be captured or proven by any algorithmic formal system like $\mathcal{F}_{QG}$. This aligns with the broader idea in the foundations of quantum gravity that there are aspects of the theory that may be fundamentally unprovable within any formal system, reinforcing the need for meta-theoretical frameworks to fully understand the working of quantum gravity.